Mullis Corp. manufactures DVDs that sell for $5.70. Fixed costs are $29,000 and variable costs are $3.70 per unit. Mullis can buy a newer production machine that will increase fixed costs by $4,350 per year, but will decrease variable costs by $0.30 per unit. What effect would the purchase of the new machine have on Mullis' break-even point in units?

8,772 unit decrease.

2,175 unit increase.

No effect.

1,891 unit decrease.

1,891 unit increase.
Answer #1

Current:

Contribution margin=Sales-Variable cost  

=5.7-3.7=$2 per unit

Breakeven=Fixed expenses/Contribution margin

=29000/2=14500 units

Now:

New variable cost=(3.7-0.3)=$3.4 per unit

Contribution margin=Sales-Variable cost  

=(5.7-3.4)=$2.3 per unit

New fixed cost=29000+4350=$33350

Hence new breakeven=33350/2.3

=14500 units

Hence there is No effect.